OpenStudy (jokersmart):

which number is irrational? A. 0.27 B. 1/9 C. 12 D. 64/4

OpenStudy (anonymous):

Well i believe that we can cancel out B and D because fractions can not be irrational This will help you decide All numbers that are not rational are considered irrational. An irrational number can be written as a decimal, but not as a fraction. An irrational number has endless non-repeating digits to the right of the decimal point.

OpenStudy (jokersmart):

\[A. 0.270 \] \[B. \frac{ 1 }{ 9 }\] \[C. \sqrt{12}\] \[D. \frac{ 64 }{ 4 }\]

OpenStudy (anonymous):

I want you to convert all the answers into decimals if you can do so.

OpenStudy (jokersmart):

Is it A?

OpenStudy (anonymous):

no it is not A because it is not repeating

OpenStudy (anonymous):

I want you to divide 26 by 3 and tell me what you get

OpenStudy (jokersmart):

8.666666666666667. There's a lot of 6's.

OpenStudy (anonymous):

that is a example of a irrational number, the decimal goes on for infinity. Your answer when in decimal form will look like this.

OpenStudy (anonymous):

So you should convert all of the answers into decimals then compare and find the irrational number